# AI-Enhanced Wi-Fi Sensing Through Single Transceiver Pair

### v1

@@ Abstract
The advancement of next-generation Wi-Fi technology heavily relies on sensing capabilities, which play a pivotal role in enabling sophisticated applications. In response to the growing demand for large-scale deployments, contemporary Wi-Fi sensing systems strive to achieve high-precision perception while maintaining minimal bandwidth consumption and antenna count requirements. Remarkably, various AI-driven perception technologies have demonstrated the ability to surpass the traditional resolution limitations imposed by radar theory. However, the theoretical underpinnings of this phenomenon have not been thoroughly investigated in existing research. In this study, we found that under hardware-constrained conditions, the performance gains brought by AI to Wi-Fi sensing systems primarily originate from two aspects: prior information and temporal correlation. Prior information enables the AI to generate plausible details based on vague input, while temporal correlation helps reduce the upper bound of sensing error. We developed an AI-based Wi-Fi sensing system using a single transceiver pair and designed experiments focusing on human pose estimation and indoor localization to validate the theoretical claims. The results confirm the performance gains contributed by temporal correlation and prior information.
@@ I Introduction
Some studies have proposed device-free Wi-Fi sensing technology suitable for large-scale applications. For instance, Widar2.0 [9] requires at least one antenna transmitter and three antenna receivers to facilitate positioning; however, its limited system resolution hampers more precise sensing applications such as human pose estimation (HPE) or breathing detection.
@@
In this paper, we first attempt to provide a theoretical explanation for the performance gains achieved by AI and then establish an Ai-based Wi-Fi sensing system using the Linux 802.11n CSI Tool [17] to validate the proposed theoretical claims. The main contributions of this work are as follows:
@@
  • We constructed an AI-based Wi-Fi sensing system using only a single transmitter-receiver pair. The system achieves an average localization error of $0.6124\text{m}$ and an average HPE error of $0.2189\text{m}$. Through experiments, we have confirmed the performance improvement brought by temporal correlation. Furthermore, we have observed indirect evidence of AI leveraging prior information.
The remainder of this paper is organized as follows. Section II presents the theoretical analysis of the performance gains achieved through AI. Section III introduces the design and implementation of the proposed Wi-Fi sensing system. Section IV details the experimental setup and results. Finally, Section V concludes the paper.
@@ II-C Performance Gain Brought by Temporal Correlation
In the presense of temporal correlation, $P_{\text{e}}$ can be represented as
@@
Figure 4: Schematic diagram of the constructed AI-based Wi-Fi sensing system, which is designed to validate the proposed theoretical claims through two tasks: human pose estimation and indoor localization.
III System Design
To validate our proposed hypothesis and theorem, we constructed a Wi-Fi based HPE system. In this section, we provide a detailed description of the system’s design.
III-A AI-based Wi-Fi Sensing System
The constructed AI-based Wi-Fi sensing system is illustrated in Fig. 4. In this system, human pose and location are estimated based on Channel State Information (CSI) collected from Wi-Fi signals. First, CSI data is acquired in an indoor environment using the Linux 802.11n CSI Tool [17]. The raw data then undergoes preprocessing to remove static components and correct phase distortions. The processed data is subsequently fed into a neural network. We validate the proposed theoretical claims through two tasks: HPE and localization. The network architecture explicitly separates localization and pose estimation. The location of a person is defined as the center of the torso, while the pose is represented by the coordinates of each joint relative to this location.
@@ III-C CSI Preprocessing
III-D Network Design
The architecture of the proposed network (Fig. 5) consists of an encoder for feature extraction, followed by a neck for intermediate processing, and a head for final pose estimation.
@@ IV-A CSI Collection
Figure 7: Example of temporal correlation gain: It can be observed that the pose estimation results become progressively more accurate over time.
@@ IV-D Experimental Results
We use the example in Fig. 7 to illustrate how temporal correlation reduces estimation errors by narrowing down the number of plausible estimation outcomes. In the figure, the blue skeleton represent the ground truth, while the red skeleton indicate the estimation results. At timestamp 1, the model infers that the subject is raising his hand, and the estimated orientation deviates significantly from the ground truth. By timestamp 2, the estimated pose shows higher similarity to the ground truth, but the orientation remains inaccurate. By timestamp 6, both the pose and orientation converge closely to the ground truth.
@@ V Conclusions
This paper theoretically analyzes and identifies two primary sources of performance gains that AI brings to Wi-Fi sensing under hardware-constrained conditions, with experimental validation of the proposed theoretical claims. We argue that the gains of AI primarily stem from the effective utilization of prior information and temporal correlation. The prior information enables AI to generate relatively detailed sensing results based on a vague perception of the target’s state by leveraging the structural priors learned during training, without the need to perceive every detail of the target meticulously. This allows AI-based sensing to surpass the system’s radar aperture. Meanwhile, temporal correlation enables AI to reduce the space of plausible sensing outcomes by leveraging the correlation across sequential data, thereby improving sensing accuracy.
We developed an AI-based Wi-Fi sensing system utilizing a single transceiver pair and validated the proposed theoretical claims through two tasks: indoor localization and human pose estimation. Experimental results confirm the performance gains brought by temporal correlation and provide indirect evidence for the benefits of prior information. Furthermore, we found that to fully leverage temporal correlation, the AI model must first undergo sufficient pre-training under a non-temporal setting to adequately learn prior structural knowledge, before being fine-tuned with temporal awareness. Otherwise, direct training tends to lead the model to converge to suboptimal local minima. The implemented system ultimately achieved an average localization error of $0.6124\text{ m}$ and an average pose estimation error of $0.2189\text{ m}$.

### v2

# AI-Enhanced Real-Time Wi-Fi Sensing Through Single Transceiver Pair

@@ Abstract
The advancement of next-generation Wi-Fi technology heavily relies on sensing capabilities, which play a pivotal role in enabling sophisticated applications. In response to the growing demand for large-scale deployments, contemporary Wi-Fi sensing systems strive to achieve high-precision perception while maintaining minimal bandwidth consumption and antenna count requirements. Remarkably, various AI-driven perception technologies have demonstrated the ability to surpass the traditional resolution limitations imposed by radar theory. However, the theoretical underpinnings of this phenomenon have not been thoroughly investigated in existing research. In this study, we found that under hardware-constrained conditions, the performance gains brought by AI to Wi-Fi sensing systems primarily originate from two aspects: prior information and temporal correlation. Prior information enables the AI to generate plausible details based on vague input, while temporal correlation helps reduce the upper bound of sensing error. Building on these insights, we developed a real-time, AI-based Wi-Fi sensing and visualization system using a single transceiver pair, and designed experiments focusing on human pose estimation and indoor localization. The system operates in real time on commodity hardware, and experimental results confirm our theoretical findings.
@@ I Introduction
Some studies have proposed device-free Wi-Fi sensing technology suitable for large-scale applications. For instance, Widar2.0 [9] requires at least a one antenna transmitter and a three antenna receiver to facilitate positioning; however, its limited system resolution hampers more precise sensing applications such as human pose estimation (HPE) or breathing detection.
@@
In this paper, we first propose a theoretical explanation for the performance gains achieved by AI. Then, based on our theoretical insight, we develop an AI-based real-time Wi-Fi sensing and visualization system and conduct experiments to validate the proposed theoretical explanation. The main contributions of this work are as follows:
@@
  • Based on our theoretical findings, we constructed an AI-based Wi-Fi sensing system using only a single transmitter-receiver pair. The system achieves an average localization error of $0.6124\text{ m}$ and an average HPE error of $0.2189\text{ m}$. Through experiments, we have confirmed the performance improvement brought by temporal correlation. Furthermore, we have observed indirect evidence of AI leveraging prior information.
  • We further optimized the data processing pipeline, thus enabling the Wi-Fi sensing system to operate in real time with simultaneous result visualization. The system achieves a frame rate of no less than $42\text{ fps}$ on commodity hardware.
The remainder of this paper is organized as follows. Section II presents the theoretical analysis of the performance gains achieved through AI. Section III introduces the design and implementation of the proposed Wi-Fi sensing system. Section IV details the experimental setup and results. Section presents the architecture of the real-time sensing and visualization system. Finally, Section VI concludes the paper.
@@ II-C Performance Gain Brought by Temporal Correlation
In the presence of temporal correlation, $P_{\text{e}}$ can be represented as
@@
Figure 4: Schematic diagram of the AI-based real-time Wi-Fi sensing system.
III AI-based Real-time Wi-Fi Sensing and Visualization System
Building on our theoretical insights, we developed an AI-based real-time Wi-Fi sensing and visualization system. In this section, we provide a detailed description of the system’s design.
III-A System Overview
The proposed AI-based real-time Wi-Fi sensing and visualization system is illustrated in Fig. 4, which estimates human pose and location from Channel State Information (CSI) extracted from Wi-Fi signals. First, CSI data is collected in an indoor environment using the Linux 802.11n CSI Tool [18]. The raw data then undergoes preprocessing to remove static components and correct phase distortions. Following this, the processed data is fed into a neural network designed based on our theory, which outputs localization and human pose estimation (HPE) results. Finally, the results are plotted and displayed in real time. Note that the location of a person is defined as the center of the torso, while the pose is represented by the coordinates of each joint relative to this location.
@@ III-C CSI Preprocessing
III-D Network Design
The architecture of the proposed network (Fig. 5) consists of an encoder for feature extraction and data compression, followed by a neck for intermediate processing, and a head for final pose estimation.
@@ III-E Loss Function
III-F Real-Time Data Processing Design
To meet real-time requirements, the following three issues must be addressed:
  • Receiver system limitation. The Linux 802.11n CSI Tool in the Wi-Fi sensing system operates on Ubuntu 14.04, which lacks support for running neural networks.
  • Inefficient reading of numerous small files. The CSI acquisition is implemented using the Linux 802.11n CSI Tool, which is written in C, whereas neural networks typically rely on Python. Thus, CSI data cannot be passed directly via memory; it must first be written to disk and then read by Python. Although CSI data is collected and saved at a rate of 1000 fps, reading it from the hard disk (HDD or SSD) is far slower, thus severely limiting the subsequent data processing frequency.
  • High overhead of scanning filenames. Reading CSI with Python requires scanning the entire directory to obtain all filenames before loading the oldest unread file to maintain sequential order. This repeated directory scanning further slows down the process.
Since the neural network structure is lightweight enough, its computational overhead is not the bottleneck for real-time performance. To resolve the above issues, we built a real-time sensing and visualization system illustrated in Fig. 6.
Figure 6: Data processing flowchart of the real-time sensing and visualization system. The receiver streams the captured CSI data in real time to a host computer. All subsequent steps — including data reading, processing, and result visualization — are integrated into a processor on the host computer.
To address the limitations of the receiver system, the receiver streams the captured CSI data in real time to a host computer. The host computer is a commodity portable personal computer, an HP OMEN 8 Pro, equipped with an NVIDIA GeForce RTX 3060 graphics card. All subsequent steps — including data reading, processing, and result visualization — are integrated into a processor on the host computer. To eliminate the overhead of scanning file names, we set up an inotify-based watcher. Whenever a new file is created, inotify generates a file update event, from which the file name can be directly obtained without scanning the entire directory. To address the inefficiency of reading a large number of small files, we adopted a multi-process and multi-thread parallel reading and processing scheme and employed queues to manage data flow between different processes. Data is pushed to the tail of the queue upon arrival, and the system pulls (and removes) data from the head of the queue when needed. This push/pull mechanism ensures that the order of data processing remains consistent with the original sequence in which the receiver collects the CSI data.
@@ IV-D Experimental Results
Figure 8: Example of temporal correlation gain, where the blue skeletons represent the ground truth, while the red skeletons indicate the estimation results. It can be observed that the pose estimation results become progressively more accurate over time.
We use the example in Fig. 8 to illustrate how temporal correlation reduces estimation errors by narrowing down the number of plausible estimation outcomes. In the figure, the blue skeletons represent the ground truth, while the red skeletons indicate the estimation results. At timestamp 1, the model infers that the subject is raising his hand, and the estimated orientation deviates significantly from the ground truth. By timestamp 2, the estimated pose shows higher similarity to the ground truth, but the orientation remains inaccurate. By timestamp 6, both the pose and orientation converge closely to the ground truth.
@@ V Conclusions
This paper theoretically analyzes and identifies two primary sources of performance gains that AI brings to Wi-Fi sensing under hardware-constrained conditions. Building on our theoretical findings, we develop an AI-based real-time Wi-Fi sensing and visualization system. We argue that the gains of AI primarily stem from the effective utilization of prior information and temporal correlation. The prior information enables AI to generate relatively detailed sensing results based on a vague perception of the target’s state by leveraging the structural priors learned during training, without the need to perceive every detail of the target meticulously. This allows AI-based sensing to surpass the system’s radar aperture. Meanwhile, temporal correlation enables AI to reduce the space of plausible sensing outcomes by leveraging the correlation across sequential data, thereby improving sensing accuracy.
The AI-based real-time Wi-Fi sensing and visualization system performs indoor localization and human pose estimation using only a single transceiver pair. The system achieved an average localization error of $0.6124\text{ m}$, an average pose estimation error of $0.2189\text{ m}$, and an average frame rate of at least $42\text{ fps}$ on commodity hardware. Experimental results confirm the performance gains brought by temporal correlation and provide indirect evidence for the benefits of prior information. Furthermore, we found that to fully leverage temporal correlation, the AI model must first undergo sufficient pre-training under a non-temporal setting to adequately learn prior structural knowledge, before being fine-tuned with temporal awareness. Otherwise, direct training tends to lead the model to converge to suboptimal local minima.